OpenStudy (anonymous):

For the system of linear equations below, use graphing to determine in which quadrant the solution lies. 3x+4y=-12 -7x+5y=-35 Quadrant I Quadrant II Quadrant III Quadrant IV

OpenStudy (zpupster):

lets eliminate on of the variables. we multiply the first equation by 7 and the 2nd eqaution by 3 getting 21x + 28y=-84 and -21x + 15y= -105 add them together ----------------- 43y =-189 y=-4.395 now you know y solve for x and plot at desmos.com or just look at the signs of your answers Quadrant I (+,+) Quadrant II (-,+) Quadrant III (-,-) Quadrant IV (+,-)

OpenStudy (anonymous):

is it quadrant 1? @zpupster

OpenStudy (zpupster):

no you see the answer for y is negative so by looking at the signs i gave for (x,y) values the y is negative in which quadrants also take the answer we got for y and plug it in and then solve for x 3x+4(-4.395)=-12

OpenStudy (anonymous):

quadrant 4? @zpupster

OpenStudy (zpupster):

yes

OpenStudy (anonymous):

thank you so much! @zpupster

OpenStudy (anonymous):

could you help me with this one too? @zpupster Consider two lines that form a linear system. Line 1: slope = 2 and y-intercept of (0,-5) Line 2: slope = 4 and y-intercept of (0,8) Part 1: What type of lines would this be? Part 2: How many solutions does this system have?